OpenStudy (anonymous):

Question 2. Find all solutions of the system of equations: −2x3 − 2x4 +x5 + 2x6 = −6 2x1 + 2x2 +x3 + x4 −4x5 − x6 = 4 2x1 + 2x2 −4x5 = 2 −x1 − x2 + x4 +4x5 = −2

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

misunderstood the question....when you say all solutions do you just mean Ax=b x=(A^-1)b

OpenStudy (zzr0ck3r):

is this for linear algebra?

OpenStudy (anonymous):

yep

OpenStudy (zzr0ck3r):

are you going over isomorphism ?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

are those exponents?

OpenStudy (zzr0ck3r):

-2x^3 or sub script -2x_3?

OpenStudy (anonymous):

done gaussian, row operations, cofactor, inverse, cramer...

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

didn't realise....2x^2 for e.g.

OpenStudy (anonymous):

not 2x2

OpenStudy (zzr0ck3r):

so they are exponents?

OpenStudy (anonymous):

yep...can be put in a matrix

OpenStudy (zzr0ck3r):

\[-x^{2}\]

OpenStudy (anonymous):

−2x^3 − 2x^4 +x^5 + 2x^6 = −6 2x^1 + 2x^2 +x^3 + x^4 −4x^5 − x^6 = 4 2x^1 + 2x^2 −4x^5 = 2 −x^1 − x^2 + x^4 +4x^5 = −2

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

when the questions asks for all solutions does it just mean Ax=b x=(A^-1)b

OpenStudy (anonymous):

so the inverse multiplied by the vector they are equal to?

OpenStudy (zzr0ck3r):

yeah, but I forgot how to do this, sorry its been a while. I thought i was onto it but I guess not.

OpenStudy (anonymous):

ok thats cool

OpenStudy (anonymous):

oh yeh i think i get it...there can only be one solution

OpenStudy (anonymous):

in this matrix i think